# Addressing Polarization and Unfairness in Performative Prediction

### v1

@@ 5 Theoretical Analysis
$t=0$, initial data distribution $\mathcal{D}^{(0)}$, strength of fair mechanism $\rho$, initial model parameter $\boldsymbol{\theta}^{(0)}$, stopping criteria $\tau$
@@
Next, we show that by replacing update $\boldsymbol{\theta}^{(t)}=\operatorname{argmin}_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})$ in Eqn. (3) with the fair update $\boldsymbol{\theta}^{(t)}=\operatorname{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ we proposed in Section 4.3, Fair-PS solutions (Definition 4.1) exist under certain conditions and the resulting iterative algorithms (see Algorithm 1) can converge to such Fair-PS solutions. We assume conditions in Lemma 2.4 hold in this section and there exists a unique PS solution in the original SDPP problem.
We first define a parameter $\tilde{\beta}$ which will be frequently used in the theorems introduced below.
@@
where ¯$\overline{\ell}:=\sup_{\boldsymbol{\theta},Z}\ell(\boldsymbol{\theta};Z)$. We can identify conditions under which a unique Fair-PS solution exists.
@@
Although the choice of repeated deployment schema does not affect the Fair-PS solution $(\mathcal{D}^{\text{PS}}_{\textsf{fair}},\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}})$, it influences the convergence rate of the iterative algorithms as stated in Theorem 5.2.
@@
Under conventional deployment schema, Algorithm 1 converges to the Fair-PS solution at a linear rate if $\epsilon(1+\tilde{\beta}/\gamma)<1$; under $k$-delayed deployment schema, Algorithm 1 converges to the Fair-PS solution at a linear rate for any $k$ if $\epsilon(1+\tilde{\beta}/\gamma)<1-\epsilon$; Particularly, under delayed deployment schema when $r=\log^{-1}\left(\frac{1}{\epsilon}\right)\log\left(\frac{\mathcal{W}_{1}(\mathcal{D}^{(0)},\mathcal{D}^{(1)})}{\delta}\right)$, Algorithm 1 converges to a radius $\delta$ of the Fair-PS solution (i.e., $\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}\|_{2}\leq\delta$ and $\mathcal{W}_{1}(\mathcal{D}(\boldsymbol{\theta}^{(t)}),\mathcal{D}^{\text{PS}}_{\textsf{fair}})\leq\delta$) in $\mathcal{O}(\log^{2}\frac{1}{\delta})$ steps if $\epsilon(1+\tilde{\beta}/\gamma)<1$.
In Appendix B.3, we extend the Algorithm 1 to practical scenarios when data distribution is inaccessible. We provide the algorithm fair repeated empirical risk minimization (Fair-RERM) where only limited data samples are available from $\mathcal{D}^{(t)}$ and analyze its sample complexity in Theorem B.3.
Impacts of strength of fair mechanism $\rho$. Theorem 5.2 and B.3 show that repeatedly minimizing $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ on evolving data sequence can lead the system converging to a Fair-PS solution. Note that $\rho$ controls the strength of fairness and different $\rho$ could result in different $(\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}},\mathcal{D}^{\text{PS}}_{\textsf{fair}})$. In conventional supervised learning with static data distribution $\mathcal{D}$, it is trivial to see that larger $\rho$ in $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D},\rho)$ will lead to a fairer solution [6]. However, in SDPP with model-dependent distribution shifts, the impact of $\rho$ on unfairness is less straightforward. Because both data distribution $\mathcal{D}^{\text{PS}}_{\textsf{fair}}$ and model $\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}$ depend on $\rho$, analyzing how group-wise disparity would change as $\rho$ varies can be highly non-trivial and has not been studied in PP literature (Appendix D.4 provides more discussion).
Next, we initiate the study for fair mechanism in Eqn. (7) with group-level fairness penalty and rigorously show that the larger $\rho$ can lead to smaller group-wise loss disparity at the Fair-PS solution. We will focus on a special case of user retention dynamics [19, 2] with two groups $s\in\{a,b\}$, where the group distribution $\mathcal{D}_{s}$ is fixed but the fraction $p_{s}^{(t)}$ changes based on group loss $\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s})$ during repeated risk minimization process.
@@
At each round, the group that is majority also experiences lower expected loss, i.e., $\arg\max_{s\in\{a,b\}}p_{s}^{t}=\arg\min_{s\in\{a,b\}}\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s})$. Moreover, for any two models deployed on population $\mathcal{D}^{(t)}$, the model with larger loss disparity $\triangle_{\mathcal{L}}^{(t)}$ leads to higher group fraction disparity $\triangle^{(t+1)}_{p}$ at time $t+1$.
Assumption 5.3 is natural: in fields such as recognition or recommendation systems, the minority group contributing less training data suffers from higher loss and has a lower retention rate.
Theorem 5.4 (Larger $\rho$ leads to stronger fairness).
Denote Fair-PS solution $\left(\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}\left(\rho\right),\mathcal{D}^{\text{PS}}_{\textsf{fair}}\left(\rho\right)\right)$ as a function of $\rho\geq 0$. Let $\triangle^{\text{PS}}_{\textsf{fair},\mathcal{L}}(\rho)$ be the group-wise loss disparity of $\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}\left(\rho\right)$ evaluated over distribution $\mathcal{D}^{\text{PS}}_{\textsf{fair}}\left(\rho\right)$ as defined in Eqn. (4). Under Assumption 5.3, $\triangle^{\text{PS}}_{\textsf{fair},\mathcal{L}}(\rho)$ is non-increasing in $\rho$.
@@ 6 Numerical Results
This section empirically evaluates the proposed methods on synthetic and real-world data (including credit data [4, 23] and MNIST data [24]) under semi-synthesized performative shifts. We run all experiments with multiple random seeds and visualize the standard error with the shaded area. Due to page limits, more experiments are presented in Appendix C.
@@
Performative Gaussian mean estimation. We generate a synthetic dataset of $10000$ samples from two groups $s\in\{a,b\}$ with initial group fractions $p_{a}^{(0)}=0.3,p_{b}^{(0)}=0.7$ and target values $y_{a}=0.3+\epsilon,y_{b}=0.7+\epsilon$, where $\epsilon\sim\mathcal{N}(0,0.05)$. We consider user retention dynamics where the group fraction $p^{(t+1)}_{s}=\frac{\mathcal{R}(s,t)}{\sum_{s^{\prime}\in\{a,b\}}\mathcal{R}(s^{\prime},t)}$ changes based on group-wise loss. Here $\mathcal{R}(s,t)=\left(1-\sum_{s^{\prime}\in\{a,b\}}p^{\min}_{s^{\prime}}\right)\times\frac{1}{2}\left(p_{s}^{t}+\frac{\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{-s}^{(t)})}{\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{-s}^{(t)})+\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s}^{(t)})}\right)+p_{s}^{\min}$ where $p_{s}^{\min}=0.02$ is the minimum group fraction that ensures the sensitivity of the transition mapping T is bounded, and $-s=\{a,b\}\setminus s$. We perform the mean estimation task and train multiple linear regression models using RERM and Fair-RERM, including regularization methods with group-level penalty (Fair-RERM-GLP), sample-level penalty (Fair-RERM-SLP) and sample re-weighting method (Fair-RERM-RW). We perform $30$ rounds of empirical risk minimization on $7$ different random seeds. Fig. 2(a) illustrates the evolution of group-wise loss disparity $\triangle_{\mathcal{L}}^{(t)}$ and Fig. 2(b) shows the evolution of the minority group fraction (i.e., the group with the lower fraction at the stable point). The blue lines in both figures demonstrate that the PS solution of the regular RERM has the highest unfairness and the most polarized group fractions, while the Fair-PS solution is fairer and the higher $\rho$ results in stronger fairness, implying Fair-RERM is effective in mitigating polarization effects and unfairness. We also note that among all three methods, the sample re-weighting method (Fair-RERM-RW) seems to yield better fairness control than others under this setting.
(a) Dynamics of group-wise loss disparity.
(b) Dynamics of the minority group fraction.
@@
Figure 2: Dynamics of unfairness under Synthetic data when $\rho\in\{0,0.3,0.7\}$ ($\rho=0$ is RERM): Fair-RERM-RW (left plot), Fair-RERM-GLP (middle plot), Fair-RERM-SLP (right plot)
@@
Credit data retention dynamics with strategic behaviors. We use the Give me some credit data [23] consisting of features $X\in\mathbb{R}^{10}$ to measure individuals’ creditworthiness $Y\in\{0,1\}$ which has been widely used in performative prediction and strategic classification literature [4, 15]. We preprocessed the data similarly to Perdomo et al. [4] and divided individuals into two groups $s\in\{a,b\}$ based on the age attribute. Next, we assume there is a newly established credit rating agency performing RERM and Fair-RERM with logistic classification models to predict individuals’ creditworthiness. Similar to the previous experiment, we assume the group-wise loss in round $t$ will affect the group fraction at $t+1$. Meanwhile, we assume there is a subset of features $X_{s}\in X$ which individuals can change strategically to $X_{s}^{\prime}$ based on the current model parameters $\theta$. Specifically, $X_{s}^{\prime}=X_{s}-\epsilon\cdot\theta_{s}$, where $\theta_{s}$ is the subset of $\theta$ with respect to $X_{s}$ and $\epsilon=0.1$. With the dynamics involving both retention and strategic behaviors, we can visualize the dynamics of unfairness under RERM and Fair-RERM in Fig. 3. All results demonstrate the effectiveness of our methods. The regularization method with sample level penalty (Fair-RERM-SLP) seems to be more effective under this setting.
@@
(a) Dynamics of group-wise loss disparity.
(b) Dynamics of the minority group fraction.
@@
Figure 3: Dynamics of unfairness under Credit data when $\rho\in\{0,1.0,5.0\}$ ($\rho=0$ is RERM): Fair-RERM-RW (left plot), Fair-RERM-GLP (middle plot), Fair-RERM-SLP (right plot)
Performative MNIST classification. We next consider MNIST data [24] consisting of handwritten digits images. We randomly select 1000 images as the training set and divide them into 2 groups, where the first group contains all images with digits from 00 to $4$ and the second includes digits from $5$ to $9$. We consider the same group fraction dynamics as before with $p_{s}^{\min}=0.1$ for both groups. Using a 2-layer MLP (the loss function violates conditions in Lemma 2.4) for risk minimization, we perform RERM and Fair-RERM for $50$ rounds. Fig. 4(a) and 4(b) verify that Fair-RERM can still improve fairness at the PS solution. Both plots demonstrate that unfairness is the largest without fairness interventions and Fair-RERM is more effective with larger $\rho$. Unlike the experiments on synthetic and credit data, this set of experiments does not have a convergence guarantee because conditions for theorems are violated under 2-layer MLP, e.g., the black and red lines ($\rho=0.3$) of Fair-RERM-RW and Fair-RERM-GLP are unstable. However, the results suggest that our fairness mechanisms may still be useful in practice with deep learning models.
(a) Dynamics of group-wise loss disparity.
(b) Dynamics of the minority group fraction.
Figure 4: Dynamics of unfairness under MNIST data when $\rho\in\{0,0.1,0.3\}$ ($\rho=0$ is RERM): Fair-RERM-RW (left plot), Fair-RERM-GLP (middle plot), Fair-RERM-SLP (right plot)
7 Conclusions & Societal Impacts

### v2

@@ 5 Theoretical analysis
0:$t=0$, initial data distribution $\mathcal{D}^{(0)}$, strength of fair mechanism $\rho$, initial model parameter $\boldsymbol{\theta}^{(0)}$, stopping criteria $\tau$
@@
In this section, we will show that by replacing $\boldsymbol{\theta}^{(t)}=\operatorname{argmin}_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})$ in Eqn. (3) with the fair update $\boldsymbol{\theta}^{(t)}=\operatorname{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ we proposed in Section 4.3, Fair-PS solutions (Definition 4.1) exist under certain conditions and the resulting iterative algorithms (Algorithm 1) can converge to such Fair-PS solutions. We assume conditions in Lemma 2.4 hold in this section and there exists a unique PS solution in the original SDPP problem. We first define a parameter $\tilde{\beta}$ which will be frequently used in the theorems introduced below.
@@
where $\overline{\ell}:=\sup_{\boldsymbol{\theta},Z}\|\ell(\boldsymbol{\theta};Z)\|$ and $\tilde{\ell}:=\sup_{\boldsymbol{\theta},Z}\{\|\nabla\ell_{\boldsymbol{\theta}}(\boldsymbol{\theta};Z)\|,\|\nabla\ell_{Z}(\boldsymbol{\theta};Z)\|\}$. We can identify conditions under which a unique Fair-PS solution exists.
@@
Although the choice of repeated deployment schema does not affect the Fair-PS solution $(\mathcal{D}^{\text{PS}}_{\textsf{fair}},\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}})$, it influences the convergence rate of the iterative algorithms in Theorem 5.2.
@@
Algorithm 1 converges to a Fair-PS solution under the following deployment schemas: (i) under conventional deployment schema, it converges to the Fair-PS solution at a linear rate if $\epsilon(1+\tilde{\beta}/\gamma)<1$; (ii) under $k$-delayed deployment schema, it converges to the Fair-PS solution at a linear rate for any $k$ if $\epsilon(1+\tilde{\beta}/\gamma)<1-\epsilon$; (iii) under delayed deployment schema when $r=\log^{-1}\left(\frac{1}{\epsilon}\right)\log\left(\frac{\mathcal{W}_{1}(\mathcal{D}^{(0)},\mathcal{D}^{(1)})}{\delta}\right)$, it converges to within a radius $\delta$ of the Fair-PS solution (i.e., $\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}\|_{2}\leq\delta$ and $\mathcal{W}_{1}(\mathcal{D}(\boldsymbol{\theta}^{(t)}),\mathcal{D}^{\text{PS}}_{\textsf{fair}})\leq\delta$) in $\mathcal{O}(\log^{2}\frac{1}{\delta})$ steps if $\epsilon(1+\tilde{\beta}/\gamma)<1$.
In App. B.3, we extend Algorithm 1 to fair repeated empirical risk minimization (Fair-RERM). Although the theorems are for fairness mechanisms in Section 4, the proofs can be easily extended to a general class of $\mathcal{L}_{\textsf{fair}}$ with convex and smooth fairness penalty terms (details in App. B.4).
Fairness guarantee.
Thm. 5.2 and App. B.3 show that repeatedly minimizing $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ on evolving data sequence can lead the system converging to a Fair-PS solution. Note that $\rho$ controls the strength of fairness and different $\rho$ could result in different $(\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}},\mathcal{D}^{\text{PS}}_{\textsf{fair}})$. In conventional supervised learning with static data distribution $\mathcal{D}$, it is trivial to see that larger $\rho$ will lead to a fairer solution (martinez2020minimax). However, in SDPP settings, the impact of $\rho$ on unfairness is less straightforward. Since both data distribution $\mathcal{D}^{\text{PS}}_{\textsf{fair}}$ and model $\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}$ depend on $\rho$, analyzing how loss disparity would change as $\rho$ varies can be highly complex. However, we manage to study fair mechanisms in Eqn. (7) with group-level fairness penalty to prove and quantify the fairness guarantee at the Fair-PS solution. We focus on a special case of user retention dynamics (hashimoto2018fairness; Zhang_2019_Retention) with two groups $s\in\{a,b\}$, where the group distribution $\mathcal{D}_{s}$ is fixed but the fraction $p_{s}^{(t)}$ changes based on group loss $\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s})$ during repeated risk minimization process.
@@
The majority group always experiences lower expected loss, i.e., $\arg\max_{s\in\{a,b\}}p_{s}^{t}=\arg\min_{s\in\{a,b\}}\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s})$. For two models deployed on population $\mathcal{D}^{(t)}$, the model with larger loss disparity $\triangle_{\mathcal{L}}^{(t)}$ leads to higher group fraction disparity $\triangle^{(t+1)}_{p}$ at time $t+1$.
Assumption 5.3 is natural: in applications such as recommendation systems, the minority group often suffers from higher loss and has a lower retention rate. Denote Fair-PS solution $\left(\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}\left(\rho\right),\mathcal{D}^{\text{PS}}_{\textsf{fair}}\left(\rho\right)\right)$ as a function of $\rho\geq 0$ and the original PS solution as $(\boldsymbol{\theta}^{\text{PS}},\mathcal{D}^{\text{PS}})$. Let $\triangle^{\text{PS}}_{\textsf{fair},\mathcal{L}}(\rho)$ be the group loss disparity of $\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}\left(\rho\right)$ evaluated on distribution $\mathcal{D}^{\text{PS}}_{\textsf{fair}}\left(\rho\right)$ and $\Delta_{\mathcal{L}}^{\text{PS}}$ be the group loss disparity of $\boldsymbol{\theta}^{\text{PS}}$ evaluated at $\mathcal{D}^{\text{PS}}$. We can first prove that a larger $\rho$ leads to stronger fairness in Thm. 5.4.
Theorem 5.4.
Under Assumption 5.3, $\triangle^{\text{PS}}_{\textsf{fair},\mathcal{L}}(\rho)$ is non-increasing in $\rho$.
Furthermore, we can quantify the fairness improvement.
Theorem 5.5.
Denote $p_{s}^{\text{PS}}$ as the fraction of group $s$ at $\mathcal{D}^{\text{PS}}$ under retention dynamics. Assume $\mathcal{L}(\boldsymbol{\theta};\mathcal{D})$ is twice continuously differentiable. For sufficiently small $\rho>0$, we have
$\Delta_{\mathcal{L}}^{\text{PS}}-\Delta_{\mathcal{L},\textsf{fair}}^{\text{PS}}(\rho)=2\rho\,p_{a}^{\text{PS}}p_{b}^{\text{PS}}\,\Delta_{\mathcal{L}}^{\text{PS}}\cdot\boldsymbol{v}_{\textsf{fair}}^{\top}H^{-1}\boldsymbol{v}_{\textsf{fair}}+\mathcal{O}(\rho^{2}),$
where $\boldsymbol{v}_{\textsf{fair}}=\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{a})-\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{b})$, and $H=\nabla_{\boldsymbol{\theta}}^{2}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}^{\text{PS}}).$
Since $H$ is positive definite due to the strong convexity of $\mathcal{L}_{\textsf{fair}}$ (Lemma D.4), Thm. 5.5 reveals the fairness improvement is positive when $\rho$ is sufficiently small even without Assumption 5.3.
@@ 6 Numerical results
This section empirically evaluates the proposed methods on synthetic and real data (including credit data (perdomo_performative_2021; creditdata) and MNIST (deng2012mnist)) under semi-synthesized performative shifts. We run all experiments with multiple random seeds and visualize the standard errors.
Performative Gaussian mean estimation.
@@
We generate a synthetic dataset of $10000$ samples from $s\in\{a,b\}$ with initial group fractions $p_{a}^{(0)}=0.3,p_{b}^{(0)}=0.7$ and target values $y_{a}=0.3+\epsilon,y_{b}=0.7+\epsilon$, where $\epsilon\sim\mathcal{N}(0,0.05)$. At each round, the decision-maker estimates the mean of the current distribution as $\boldsymbol{\theta}^{(t)}$, and the loss function is the mean squared error with $l_{2}$ regularization. We consider user retention dynamics similar to hashimoto2018fairness where $p^{(t+1)}_{s}=\frac{\mathcal{R}(s,t)}{\sum_{s^{\prime}\in\{a,b\}}\mathcal{R}(s^{\prime},t)}$ changes based on group-wise loss. Here $\mathcal{R}(s,t)=\left(1-\sum_{s^{\prime}\in\{a,b\}}p^{\min}_{s^{\prime}}\right)\times\frac{1}{2}\left(p_{s}^{t}+\frac{\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{-s}^{(t)})}{\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{-s}^{(t)})+\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s}^{(t)})}\right)+p_{s}^{\min}$ where $p_{s}^{\min}=0.02$ is the minimum group fraction, and $s=\{a,b\}\setminus s$. We perform the mean estimation task empirically to train multiple linear regression models and compare RERM with Fair-RERM, including regularization methods with group-level penalty (Fair-RERM-GLP), sample-level penalty (Fair-RERM-SLP) and sample re-weighting method (Fair-RERM-RW). We perform $30$ rounds of empirical risk minimization on $7$ different random seeds. Fig. 2(a) illustrates the evolution of group-wise loss disparity $\triangle_{\mathcal{L}}^{(t)}$, where higher fairness improvement is achieved with higher $\rho$. The sample re-weighting method (Fair-RERM-RW) seems to yield better fairness control than others in this setting. Fig. 2(b) shows the tradeoff between fairness (group-wise loss disparity) and the global loss. The “stars” are produced by adjusting $\rho$ and demonstrate a “Pareto-optimal” surface of each fairness mechanism.
(a) Dynamics of loss disparity for Performative Gaussian mean estimation when $\rho\in\{0,1.0,2.0\}$. When $\rho=0$, the policy reduces to standard RERM.
(b) Tradeoff between fairness and the global loss at the PS solution. “Stars” are produced by varying $\rho$.
@@
Figure 2: Results on Gaussian data: Fair-RERM-RW (left), Fair-RERM-GLP (middle), Fair-RERM-SLP (right).
Credit data retention dynamics with strategic behaviors.
@@
We use the Give Me Some Credit data (creditdata) consisting of features $X\in\mathbb{R}^{10}$ to measure individuals’ creditworthiness $Y\in\{0,1\}$ (perdomo_performative_2021; hu2022achieving). We preprocessed the data similarly to perdomo_performative_2021 and divided individuals into two groups $s\in\{a,b\}$ based on the age attribute. Next, we assume there is a newly established credit rating agency comparing RERM with Fair-RERM with logistic classification models to predict individuals’ creditworthiness. Similarly, we assume the group-wise loss in round $t$ affects the group fraction at $t+1$. Meanwhile, we assume there is a subset of features $X_{s}\in X$ which individuals can change strategically to $X_{s}^{\prime}$ based on the current model parameters $\theta$. Specifically, $X_{s}^{\prime}=X_{s}-\epsilon\cdot\theta_{s}$, where $\theta_{s}$ is the subset of $\theta$ with respect to $X_{s}$ and $\epsilon=0.1$. With the dynamics, we can visualize the evolution of group-wise loss disparity and the tradeoff between fairness and the global loss in Fig. 3. All results demonstrate the effectiveness of our methods where Fair-RERM-SLP seems to be more effective.
@@
(a) Dynamics of group-wise loss disparity for Credit data when $\rho\in\{0,1.0,5.0\}$. When $\rho=0$, the policy is just RERM.
(b) Tradeoff between fairness and the global loss at the PS solution. “Stars” are produced by varying $\rho$.
@@
Figure 3: Results on Credit data: Fair-RERM-RW (left), Fair-RERM-GLP (middle), Fair-RERM-SLP (right).
Additional experiments in App. C.
Due to the page limit, we defer additional experiments to App. C.1. We want to highlight that we perform experiments on MNIST data (deng2012mnist) to test whether our fairness mechanisms can still be useful beyond the convex setting with a deep learning model. Remarkably, Fig. 6(a) in App. C.1 verifies that Fair-RERM can still improve fairness at the PS solution. Moreover, Fig. 6(b) in App. C.1 demonstrates that the Fair-PS solution may converge to a local stationary point better for both fairness and performative loss in non-convex PP settings. We also perform experiments on ACSIncome dataset (ding2022retiringadultnewdatasets) (App. C.3) and show how our fairness mechanisms influence Equal opportunity and Demographic Parity (App. C.7).
7 Conclusions & limitations